9*x^2+25*y^2=225
Solve for y in terms of x.

Answer by jim_thompson5910(35256)

9x^2+25y^2=225

25y^2=225-9x^2

y^2=(225-9x^2)/25

y=+-sqrt((225-9x^2)/25)

y=sqrt(225-9x^2)/5 or y=-sqrt(225-9x^2)/5